OpenStudy (anonymous):

A person who is 5ft 8in tall cast a shadow that is 12 feet long. The distance along the ground to the flagpole is 18 feet. find the height of the flagpole

OpenStudy (ddcamp):

8 inches = 2/3 of a foot, 5ft = 15/3 of a foot. The person is 17/3 of a foot tall: